ANNOUNCEMENTS ------------- * For the remainder of the week, call Stein or Emily in case you need to get hold of a SOC. * HESSI is ready to begin joint observations. HESSI target: NOAA 9825 * NRT plans for the immediate future: Today most teams will command on D66 in the morning, CDS and MDI will command on D27 in the afternoon. Friday we have D66 and D16 during the day. The Intercal goes at 15:00, and LASCO loads a new plan afterwards, before the handover. During the handover, MDI may ftp over a background load. CDS starts an FF_CAL study after the handover, together with MDI's background load uplinking (or MDI NRT). CDS will also load science plans for the weekend. Station 27 comes up starting around 22:00. Note that as soon as we get final plans for the remainder of the CELIAS/CTOF recovery, we will try to schedule that NRT pause at a convenient time during daytime. * Commanding schedule for next week: Monday - D66 until 16:30, D27 after 19 UT. Tuesday - D66 until 18:30 then D27. Wedesday - D66 until 18:30 then D27. Thursday - D/L only in the morning then D27 after 17:30 UT. Friday - D/L only until 19 UT then D16. Saturday - Command on D27 all day. Sunday - Command on D27 all day. FOT REPORT ---------- Spacecraft Status: SVM Nominal, SUMER OFF MDI, CTOF not yet in nominal operations Spacecraft Anomalies: None Accomplished Activities: VIRGO, VC2 (1600-0200) Planned Activities: VIRGO, NOM HGA Table, VC2 (1450-0800) Upcoming Operations: None Ground Anomalies: 044/1421 - TM Dropout; D66 rcvr problem (3 minutes, unrecoverable) 044/1546 - NRT Interrupted; H/W Maintenance inadvertently pulled cable 045/0837 - Intermittent lock on RCVR2 during D46 pass SOLAR STATUS ------------ 02/13 DOY 044 - 02/14 DOY 045 EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20020214.html LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/2001/Todays_obs.html Web Page for Planning: http://www.bbso.njit.edu/arm INSTRUMENT STATUS ----------------- CDS: Nominal. Polarity reversal study in the south, active region study near central meridian. Tomorrow: ICAL, studies TBD. UVCS: Nominal. Coronal hole observations and synoptic for today and tomorrow. LASCO: Nominal. C2/C3 synoptics. EIT: Nominal. 195 Angstrom CME watch and synoptics. MDI: Restarted structure and closed the loop yesterday, taking some magnetograms and continuum images. Loads (w/pause in structure program) and calibrations today. Only a few pixels in the flat field needs to be patched. Hope to be back in normal science mode today. TRACE: Nominal. Continue observing AR#9825 and the other regions of the complex; using both the 1600A flarewatch/respond and white light flarewatch/respond sequences. Goal is to co-observe flares with HESSI.